OpenStudy (anonymous):

Convert 3x^2 + 3y^2 + 10 to polar form

OpenStudy (anonymous):

is that =0?

OpenStudy (anonymous):

Oops sorry. It's suppose to be 3x^2 + 3y^2 = 10

OpenStudy (anonymous):

Ah. I suspected as much. Do you know what the graph of 3x² + 3y² = 10 looks like?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Im thinking the answer would be r^2 = 10/3

OpenStudy (anonymous):

Yeah, that's correct, although I'd write it in the form \[r = \sqrt{\frac{10}{3}}\]